OpenStudy (anonymous):

A bag contains 26 letters and 9 numerals. Aubrey draws two cards from the bag. What is the probability that she draws 2 numerals?

OpenStudy (anonymous):

ok well you have 9 numerals so u gonna put 2 over 9 so it makes it a fraction so 2/9 and you cant simplify 2/9 so theirs ur answer

OpenStudy (anonymous):

My bad i forgot to put the answer choices. A.36/595 <- Is my current answer im not sure if its right though, so i'd like if someone could show me how they got their answer if im wrong B.2/35 C.13/127 D.1/14

OpenStudy (anonymous):

26+9 = 35 35C2 = 595

OpenStudy (anonymous):

35/595

OpenStudy (anonymous):

Your answer is correct. \[\frac{\dbinom{26}0 \dbinom92}{\dbinom{35}2}=\frac{36}{595}\]

OpenStudy (anonymous):

Thank you everyone for your help!